# AttributionLab: Faithfulness of Feature Attribution Under Controllable Environments

### v1

@@ B.1.1 Gradient-based
The early methods relied on a linear approximation of the neural network around the input point. Simonyan et al. (2013) assumed the model to be linear (inspired by the piecewise linearity of ReLU) and proposed using the gradient of output with respect to input as a saliency map. This work coined the term “saliency map” for attribution, and many following works used the same nomenclature. The gradient of neural networks with respect to images looked “noisy”. Therefore, there were attempts to remove the ”noise” from the images. One of these methods averages the gradient over a neighborhood around the input to smooth out the gradient. The SmoothGrad Smilkov et al. (2017) method adds Gaussian noise to the image and generates multiple samples (in a Gaussian neighborhood of the input). Then computes the gradient and averages it. Other works proposed backpropagating only positive values to remove the ”noise”. Deconvolution Zeiler & Fergus (2014) and subsequently Guided Backpropagation Springenberg et al. (2015) and Excitation BackProp Zhang et al. (2018) work based on backpropagating positive values.
@@ B.1.2 Latent features
The most famous work from this category is CAM/GradCAM Selvaraju et al. (2017); Zhou et al. (2016). The method does a weighted average of activation values of the final convolutional layer. The weights are selected based on their gradient values. Another work Khakzar et al. (2021) computes the input attribution based on the contribution of neurons to the output. This is achieved by computing the gradient of critical pathways. Another method restricts the flow of information in the latent space Schulz et al. (2020) and proposes information bottleneck attribution (IBA).
@@ B.1.3 Backpropagation of relevance
The most important works within this category are LRP Bach et al. (2015), and DeepLift Shrikumar et al. (2017). The LRP method provides a general framework for backpropagating the relevance to input while satisfying a conservation property (similar to Kirchoff’s law of electricity). The relevance entering a neuron equals the relevance coming out. There are several rules of backpropagation within the LRP framework, each with its own properties. DeepLift proposes a chain rule for propagating output differences to the input. An improved version of DeepLift is the DeepSHAP Lundberg & Lee (2017) method, which we use in the experiments.
@@ B.1.4 Perturbation-based methods
Another category of methods performs attribution by perturbing the input. The most straightforward way is to occlude/mask pixels (or patches of pixels) within the image and measure the output difference Zeiler & Fergus (2014); Ancona et al. (2018). The output difference reflects the importance of the removed feature. It is also possible to search for a mask on the input image. For instance, searching for the smallest region within the image which preserves the output Fong et al. (2019).
@@ B.1.5 Shapley value
The Shapley value Shapley (1953) itself is based on the occlusion of features. It is a notion from cooperative game theory for assigning the contribution of players to a game. We can consider the pixels within the image as players and the output of the network as the score of the game. The Shapley value is a unique solution that satisfies the symmetry, dummy, linearity, and completeness axioms altogether. It is of exponential complexity and, therefore, almost impossible to compute for large images and neural networks. Lundberg & Lee (2017) first proposed using the Shapley value (an approximation) for feature attribution in neural networks by introducing the SHAP and DeepSHAP (for neural networks) methods. Another method based on the Shapley value is the integrated gradients Sundararajan et al. (2017), which is the Aumann-Shapley value in the continuous domain Sundararajan & Najmi (2020).
@@ B.2.1 Alignment with human intuition
Early feature attribution works Simonyan et al. (2013); Zeiler & Fergus (2014); Zhang et al. (2018); Bach et al. (2015) evaluated different saliency methods based on how they are aligned with what we humans think is salient. For instance, if a saliency method highlights a certain object, we might conclude that the network uses this object. However, the network may be using different features than ours. In this case, we wrongly consider the attribution method as correct (known as confirmation bias). Or in another case, the network might be using the object, but the saliency method might highlight something else. How would we know that the saliency is wrong? Several methods, such as the pointing game Zhang et al. (2018), made this visual evaluation systematic by comparing the saliency values with ground truth bounding box annotations. Even though human evaluation is flawed (e.g., due to confirmation bias), it is still useful for debugging and better understanding attributions.
@@ B.2.2 Sanity checks
Another group of metrics evaluates the sanity of methods. They check if the methods possess specific properties that attribution methods must have. Adebayo et al. (2018) checks what happens to attributions when we replace the network weights with random values. It is surprisingly observed that some methods generate the same saliency values when the network is randomized. Another work Sixt et al. (2020) checks how the attribution method changes when it is applied to a different output. In the presence of multiple classes within the image, the attribution method is expected to point to the relevant features of the explained class.
@@ B.2.3 Evaluation by perturbation
The intuition behind these methods is aligned with the perturbation-based feature attribution methods. If a feature is contributing to the output, its removal should affect the output. And the more the effect, the more important the feature. Based on this intuition, Samek et al. (2016) proposed perturbing/removing pixels based on their attribution score and plotting their effect as we keep removing them. However, one issue with this approach is that the output score might be due to out-of-distribution values and not the relevance of the feature. Another method Hooker et al. (2019) tries to remedy this issue by training the neural network from scratch on the perturbed dataset. In other words, a certain percentage of each image within the dataset is perturbed based on attribution scores, and the network is trained on this new dataset. The method then computes the accuracies in different perturbation percentages.
@@ B.2.4 Axiomatic
Several axioms provide a framework to formalize the contribution of a feature Sundararajan & Najmi (2020); Lundberg & Lee (2017); Sundararajan et al. (2017). For instance, if the perturbation of a feature does not affect the output in any combination of features, then it is said to be a dummy feature. The dummy axiom demands that an attribution method should assign zero relevance to a dummy feature. Several other axioms reflect properties that we want the attributions to conform to. A combination of four properties, dummy, symmetry, linearity, and completeness, is only satisfied by the Shapley value Shapley (1953); Sundararajan et al. (2017). However, the Shapely value is not computable for large images and networks. Moreover, the proper reference value for removal must be chosen for computing the Shapley value. The complexity and the choice of reference values remain open problems. Moreover, for the existing methods, it is complex to theoretically show whether they conform to axioms as the methods are complex to analyze on neural networks. It is also shown that several axiomatic methods break the axioms in practical applications Sundararajan & Najmi (2020).
@@ B.2.5 Alignment with synthetic ground-truth
An intriguing and promising solution is proposing synthetic settings to evaluate explanations. Several approaches propose generating datasets with known correlations between features and labels Arras et al. (2022); Agarwal et al. (2023); Zhou et al. (2022). However, there is no guarantee that the network will use the same features to predict the outputs. Therefore, we will not have a ground truth of which feature is relevant to the network. This category is the focus of this work. For further information, please refer to the main text.
@@ Appendix D Implementation of Single-color-modulo Setting
Appendix E Experiments in the single-color-modulo setting
E.1 Synthetic dataset generation
@@
E.2 Hyper-parameters of attribution methods
@@
E.3 Ground-truth-based evaluation
@@ E.3.2 Experimental results
E.4 Insertion/Deletion
E.4.1 Experimental setup
@@
E.4.2 Experimental results
As illustrated in Table 3, the best performing methods are IG (or $\text{IG}^{*}$), DeepSHAP, and LRP, achieving optimal Insertion AUC and Deletion AUC. As indicated by the Spearman’s rank correlations in Tables 1 and 2, the Insertion/Deletion evaluation is highly consistent with the ground-truth-based evaluation. It is important to note that the optimal Insertion AUC in Table 3 is not $1.0$, and the optimal Deletion AUC is not $0.0$. This occurs because we remove one pixel at each step, and the best Insertion curve is obtained when all $1$s in the image receive higher attribution than all 00s. In this case, the Insertion curve is a monotone increasing straight line followed by a saturated flat line at $1.0$. Similarly, the optimal Deletion curve occurs when all the $1$s are removed before all the 00s, resulting in a monotone decreasing straight line followed by a stagnant flat line at $0.0$. Consequently, the AUC of the optimal Insertion curve is smaller than $1.0$, and the AUC of the optimal Deletion curve is greater than $0.0$. The optimal AUCs depend on the number of pixels with value $1$ and the total number of pixels.
@@ E.5.2 Experimental results
Appendix F Experiments in the multi-color-sum setting
@@
F.1 Synthetic dataset generation
@@
F.2 Hyper-parameters of attribution methods
@@
  • LIME: The default segmentation method is Quickshift Vedaldi & Soatto (2008).
@@
  • Occlusion: A sliding window of shape $(3,5,5)$ and strides $(3,3,3)$ is used, along with an all-zero baseline. Although we know the ground truth baseline, we do not use it here because the true baseline is difficult to access in real-world applications, and we aim to study the influence of an inappropriate baseline in the multi-color setting. Furthermore, the output change is measured at the layer before the softmax, as suggested by the original paper Zeiler & Fergus (2014).
@@
  • IG: The output change is measured after the softmax layer, and an all-zero baseline is employed.
@@
F.3 Ground-truth-based evaluation
F.3.1 Experimental setup
@@ F.3.2 Experimental results
F.4 Insertion/Deletion
F.4.1 Experimental setup
@@ F.4.2 Experimental results
F.5 Sensitivity-N
F.5.1 Experimental setup
@@
F.5.2 Experimental results
The Sensitivity-N curves are shown in Figure 7(c). First, we observe that the performance of the attribution methods is much less distinguishable than in the single-color-modulo case. For instance, the curves of GradCAM, GuidedBP, and Occlusion intersect with one another multiple times, making it harder to conclude which method displays superior performance. Second, DeepSHAP produces a negative correlation curve. This can be interpreted by our previous analysis that DeepSHAP often misidentifies the sign of contributing pixels. Therefore, the sum of attribution of perturbed pixels is much less aligned with the predicted probability of the perturbed image. Third, in order to rank these methods and compare the ranking with the ranking on GT-based evaluation, we simply compute the average correlation over all $N$s. As demonstrated in Table 4, the Sensitivity-N evaluation result is much less correlated with the ground-truth-based evaluation compared to that in the single-color-modulo setting.
Appendix G Experiments in the multi-color-sum setting (without Unseen Data Effect)
In our default settings for the Accumulator module within the multi-color setting, we employ a non-uniform weight initialization. In this section, we present supplementary empirical results from the multi-color-sum setting, where we implement a uniform weight initialization in the Accumulator. Please note that: (1) the weight initialization schemes do not impact the Accumulator output, ensuring that the model still achieves $100\%$ accuracy on the synthetic dataset; and (2) the color detector retains redundant channels and consequently, remains susceptible to Unseen Data Effect. In the following subsections, the evaluation settings are identical as the Appendix F, except for the number of redundant channels in the Accumulator.
@@ Appendix G Experiments in the multi-color-sum setting (without Unseen Data Effect)
G.1 Ground-truth-based evaluation
@@ G.2 Insertion/Deletion
G.3 Sensitivity-N
Figure 7(d) depicts the correlation curves.
Appendix H Broader impacts

### v2

@@ B.1.1 Gradient-based
The early methods relied on a linear approximation of the neural network around the input point. (Simonyan et al., 2013) assumed the model to be linear (inspired by the piecewise linearity of ReLU) and proposed using the gradient of output with respect to input as a saliency map. This work coined the term “saliency map” for attribution, and many following works used the same nomenclature. The gradient of neural networks with respect to images looked “noisy”. Therefore, there were attempts to remove the ”noise” from the images. One of these methods averages the gradient over a neighborhood around the input to smooth out the gradient. The SmoothGrad (Smilkov et al., 2017) method adds Gaussian noise to the image and generates multiple samples (in a Gaussian neighborhood of the input). Then computes the gradient and averages it. Other works proposed backpropagating only positive values to remove the ”noise”. Deconvolution (Zeiler & Fergus, 2014) and subsequently Guided Backpropagation (Springenberg et al., 2015) and Excitation BackProp (Zhang et al., 2018) work based on backpropagating positive values.
@@ B.1.2 Latent features
The most famous work from this category is CAM/GradCAM (Selvaraju et al., 2017; Zhou et al., 2016). The method does a weighted average of activation values of the final convolutional layer. The weights are selected based on their gradient values. Another work (Khakzar et al., 2021) computes the input attribution based on the contribution of neurons to the output. This is achieved by computing the gradient of critical pathways. Another method restricts the flow of information in the latent space (Schulz et al., 2020) and proposes information bottleneck attribution (IBA).
@@ B.1.3 Backpropagation of relevance
The most important works within this category are LRP (Bach et al., 2015), and DeepLift (Shrikumar et al., 2017). The LRP method provides a general framework for backpropagating the relevance to input while satisfying a conservation property (similar to Kirchoff’s law of electricity). The relevance entering a neuron equals the relevance coming out. There are several rules of backpropagation within the LRP framework, each with its own properties. DeepLift proposes a chain rule for propagating output differences to the input. An improved version of DeepLift is the DeepSHAP (Lundberg & Lee, 2017) method, which we use in the experiments.
@@ B.1.4 Perturbation-based methods
Another category of methods performs attribution by perturbing the input. The most straightforward way is to occlude/mask pixels (or patches of pixels) within the image and measure the output difference (Zeiler & Fergus, 2014; Ancona et al., 2018). The output difference reflects the importance of the removed feature. It is also possible to search for a mask on the input image. For instance, searching for the smallest region within the image which preserves the output (Fong et al., 2019).
@@ B.1.5 Shapley value
The Shapley value (Shapley, 1953) itself is based on the occlusion of features. It is a notion from cooperative game theory for assigning the contribution of players to a game. We can consider the pixels within the image as players and the output of the network as the score of the game. The Shapley value is a unique solution that satisfies the symmetry, dummy, linearity, and completeness axioms altogether. It is of exponential complexity and, therefore, almost impossible to compute for large images and neural networks. (Lundberg & Lee, 2017) first proposed using the Shapley value (an approximation) for feature attribution in neural networks by introducing the SHAP and DeepSHAP (for neural networks) methods. Another method based on the Shapley value is the integrated gradients (Sundararajan et al., 2017), which is the Aumann-Shapley value in the continuous domain (Sundararajan & Najmi, 2020).
@@ B.2.1 Alignment with human intuition
Early feature attribution works (Simonyan et al., 2013; Zeiler & Fergus, 2014; Zhang et al., 2018; Bach et al., 2015) evaluated different saliency methods based on how they are aligned with what we humans think is salient. For instance, if a saliency method highlights a certain object, we might conclude that the network uses this object. However, the network may be using different features than ours. In this case, we wrongly consider the attribution method as correct (known as confirmation bias). Or in another case, the network might be using the object, but the saliency method might highlight something else. How would we know that the saliency is wrong? Several methods, such as the pointing game (Zhang et al., 2018), made this visual evaluation systematic by comparing the saliency values with ground truth bounding box annotations. Even though human evaluation is flawed (e.g., due to confirmation bias), it is still useful for debugging and better understanding attributions.
@@ B.2.2 Sanity checks
Another group of metrics evaluates the sanity of methods. They check if the methods possess specific properties that attribution methods must have. (Adebayo et al., 2018) checks what happens to attributions when we replace the network weights with random values. It is surprisingly observed that some methods generate the same saliency values when the network is randomized. Another work (Sixt et al., 2020) checks how the attribution method changes when it is applied to a different output. In the presence of multiple classes within the image, the attribution method is expected to point to the relevant features of the explained class.
@@ B.2.3 Evaluation by perturbation
The intuition behind these methods is aligned with the perturbation-based feature attribution methods. If a feature is contributing to the output, its removal should affect the output. And the more the effect, the more important the feature. Based on this intuition, (Samek et al., 2016) proposed perturbing/removing pixels based on their attribution score and plotting their effect as we keep removing them. In contrast to perturbing pixels according to a specific order, Sensitivity-N (Ancona et al., 2018) randomly selects $N$ pixels to perturb, after that measuring the correlation between the change in prediction and the sum of attribution of the perturbed pixels. However, one issue with this approach is that the output score might be due to out-of-distribution values and not the relevance of the feature. Another method (Hooker et al., 2019) tries to remedy this issue by training the neural network from scratch on the perturbed dataset. In other words, a certain percentage of each image within the dataset is perturbed based on attribution scores, and the network is trained on this new dataset. The method then computes the accuracies in different perturbation percentages.
@@ B.2.4 Axiomatic
Several axioms provide a framework to formalize the contribution of a feature (Sundararajan & Najmi, 2020; Lundberg & Lee, 2017; Sundararajan et al., 2017). For instance, if the perturbation of a feature does not affect the output in any combination of features, then it is said to be a dummy feature. The dummy axiom demands that an attribution method should assign zero relevance to a dummy feature. Several other axioms reflect properties that we want the attributions to conform to. A combination of four properties, dummy, symmetry, linearity, and completeness, is only satisfied by the Shapley value (Shapley, 1953; Sundararajan et al., 2017). However, the Shapely value is not computable for large images and networks. Moreover, the proper reference value for removal must be chosen for computing the Shapley value. The complexity and the choice of reference values remain open problems. Moreover, for the existing methods, it is complex to theoretically show whether they conform to axioms as the methods are complex to analyze on neural networks. It is also shown that several axiomatic methods break the axioms in practical applications (Sundararajan & Najmi, 2020).
@@ B.2.5 Alignment with synthetic ground-truth
An intriguing and promising solution is proposing synthetic settings to evaluate explanations. Several approaches propose generating datasets with known correlations between features and labels (Arras et al., 2022; Agarwal et al., 2023; Zhou et al., 2022). However, there is no guarantee that the network will use the same features to predict the outputs. To address this problem, Khakzar et al. (2022) proposes a post-hoc solution of inverse feature generation by generating input features using the network. However, this direction faces challenges related to generated out-of-distribution input features. It is still unclear which parts of the generated features contribute and how they contribute. Therefore, we will not have a ground truth of which feature is relevant to the network. This category is the focus of this work. For further information, please refer to the main text.
@@ Appendix D Implementation of Single-color-modulo Setting
Appendix E Experiment detail for Table 1
The model summary is conducted on two models.
Ours is the designed model for the single-color-modulo setting. The model is designed to have two modules, one CNN model for pixel accumulation, and a Linear module for modulo calculation.
The CNN module comprises 10 CNN layers. The linear module has of 7 linear layers as described in the Appendix C.3. ResNet-8 model has 3 residual blocks with 2 CNN layers in each residual block, a CNN layer at the beginning, and a linear layer at the end.
Appendix F Experiments in the single-color-modulo setting
F.1 Synthetic dataset generation
@@
F.2 Hyper-parameters of attribution methods
@@
F.3 Ground-truth-based evaluation
@@ F.3.2 Experimental results
F.4 Insertion/Deletion
F.4.1 Experimental setup
@@
F.4.2 Experimental results
As illustrated in Table 4, the best performing methods are IG (or $\text{IG}^{*}$), DeepSHAP, and LRP, achieving optimal Insertion AUC and Deletion AUC. As indicated by the Spearman’s rank correlations in Tables 2 and 3, the Insertion/Deletion evaluation is highly consistent with the ground-truth-based evaluation. It is important to note that the optimal Insertion AUC in Table 4 is not $1.0$, and the optimal Deletion AUC is not $0.0$. This occurs because we remove one pixel at each step, and the best Insertion curve is obtained when all $1$s in the image receive higher attribution than all 00s. In this case, the Insertion curve is a monotone increasing straight line followed by a saturated flat line at $1.0$. Similarly, the optimal Deletion curve occurs when all the $1$s are removed before all the 00s, resulting in a monotone decreasing straight line followed by a stagnant flat line at $0.0$. Consequently, the AUC of the optimal Insertion curve is smaller than $1.0$, and the AUC of the optimal Deletion curve is greater than $0.0$. The optimal AUCs depend on the number of pixels with value $1$ and the total number of pixels.
@@ F.5.2 Experimental results
Appendix G Experiments in the multi-color-sum setting
@@
G.1 Synthetic dataset generation
@@
G.2 Hyper-parameters of attribution methods
@@
  • LIME: The default segmentation method is Quickshift (Vedaldi & Soatto, 2008).
@@
  • Occlusion: A sliding window of shape $(3,5,5)$ and strides $(3,3,3)$ is used, along with the ground truth baseline. Furthermore, the output change is measured at the layer before the softmax, as suggested by the original paper (Zeiler & Fergus, 2014).
@@
  • IG: The output change is measured after the softmax layer, and the ground truth baseline is employed.
@@
G.3 Ground-truth-based evaluation
G.3.1 Experimental setup
@@ G.3.2 Experimental results
G.4 Insertion/Deletion
G.4.1 Experimental setup
@@ G.4.2 Experimental results
G.5 Sensitivity-N
G.5.1 Experimental setup
@@
G.5.2 Experimental results
Figure 20: Sensitivity-N in the multi-color-sum setting.
The Sensitivity-N curves are shown in Figure 20. First, we observe that the performance of the attribution methods is much less distinguishable than in the single-color-modulo case. For instance, the curves of GradCAM, GuidedBP, and Occlusion intersect with one another multiple times, making it harder to conclude which method displays superior performance. Second, DeepSHAP produces a negative correlation curve. This can be interpreted by our previous analysis that DeepSHAP often misidentifies the sign of contributing pixels. Therefore, the sum of attribution of perturbed pixels is much less aligned with the predicted probability of the perturbed image. Third, in order to rank these methods and compare the ranking with the ranking on GT-based evaluation, we simply compute the average correlation over all $N$s. As demonstrated in Table 5, the Sensitivity-N evaluation result is much less correlated with the ground-truth-based evaluation compared to that in the single-color-modulo setting.
Appendix H Experiment Setting for Figure 2
All attribution maps generated in the same setting that is discussed in Appendix G.
Appendix I Experiments in the multi-color-sum setting (without Unseen Data Effect)
In our default settings for the Accumulator module within the multi-color setting, we employ a non-uniform weight initialization. In this section, we present supplementary empirical results from the multi-color-sum setting, where we implement a uniform weight initialization in the Accumulator. Please note that: (1) the weight initialization schemes do not impact the Accumulator output, ensuring that the model still achieves $100\%$ accuracy on the synthetic dataset; and (2) the color detector retains redundant channels and consequently, remains susceptible to Unseen Data Effect. In the following subsections, the evaluation settings are identical as the Appendix G, except for the number of redundant channels in the Accumulator.
@@ Appendix I Experiments in the multi-color-sum setting (without Unseen Data Effect)
Figure 21: Sensitivity-N in multi-color-sum setting without the Unseen Data Effect
I.1 Ground-truth-based evaluation
@@ I.2 Insertion/Deletion
I.3 Sensitivity-N
Figure 21 depicts the correlation curves.
Appendix J Broader impacts
@@ I.3 Sensitivity-N
Appendix K Additional study on segmentation methods of LIME
In Section 4.2, we observed that the segmentation method significantly affects LIME’s performance. This observation prompted us to explore a more advanced segmentation approach capable of effectively grouping features that function together and identifying independent feature groups. We conducted a comparative analysis of LIME attribution maps generated using Quickshift (Vedaldi & Soatto, 2008), Felzenszwalb (Felzenszwalb & Huttenlocher, 2004), and the recent segmentation approach Segment Anything Model (SAM) (Kirillov et al., 2023). Our comparison in the multi-color-sum setting, as shown in Figure 24, reveals a substantial increase in faithfulness when SAM is utilized as the segmentation method. Furthermore, we employ LIME for explaining two ImageNet-pretrained models: VGG (Simonyan & Zisserman, 2015) and ViT (Dosovitskiy et al., 2020), respectively. Visual inspection of the attribution maps on VGG16 (Figure 25) and ViT-B-16 (Figure 26), indicates that SAM provides a prior of more cohesively grouped features and significantly reduces noise in the LIME attribution maps. These attribution maps are much better aligned with human perception as well. Moreover, this finding is consistent across the CNN and ViT architectures. This study underscores that the findings in AttributionLab can assist researchers to identify strategies for enhancing existing attribution methods, and it demonstrates that these enhancements can generalize to complex real-world scenarios.
Figure 24: LIME in the multi-color-sum setting with different segmentation methods. In contrast to Quickshift (Vedaldi & Soatto, 2008) and Felzenszwalb (Felzenszwalb & Huttenlocher, 2004), SAM (Kirillov et al., 2023) offers markedly improved segmentation masks for inputs. This improvement significantly boosts the faithfulness of attribution maps, ensuring closer alignment with the ground truth attribution map.
Figure 25: LIME attribution maps for VGG16 with different segmentation methods. In line with results in the multi-color-sum setting, SAM segments input features into more compact and semantically meaningful groups. Consequently, the LIME attribution maps generated using SAM exhibit reduced noise and are visually more congruent with human perception.
Figure 26: LIME attribution maps for ViT-B-16 with different segmentation methods. These attribution maps show a significant enhancement in attribution quality when LIME incorporates SAM, and this finding generalizes to vision transformers.